OpenStudy (anonymous):

Please help me I have a test in an hour :( Simplify: (x^5)^2y^4 / (2xy)^3 I need the working out also, thanks in advance sooo much

OpenStudy (anonymous):

How do you get the answer?

OpenStudy (loser66):

\((x^5)^2 = x^{10}\) ok? I break it to smallest term

OpenStudy (loser66):

so, the numerator is \(x^{10} y^4\) got this part?

OpenStudy (anonymous):

Ok...

OpenStudy (loser66):

for denominator, you distribute the exponent to EACH of the term so that \(\Large (2xy)^3 = 2^3*x^3*y^3= 8x^3y^3\) Got this part?

OpenStudy (anonymous):

Yes

OpenStudy (loser66):

Now, combine both numerator and denominator \[\Large \dfrac{x^{10}y^4}{8x^3y^3}= \dfrac{x^7y}{8}\]

OpenStudy (loser66):

Got it?

OpenStudy (anonymous):

Yes thanks a ton :D

OpenStudy (loser66):

np